# Notes on Deterministic and Stochastic Approaches in Electromagnetic Information Theory

### v1

@@ 4 Spatially incoherent homogeneous linear source
Now, let us consider the EIT stochastic approach.
The mutual information is given by Eq. (19), with $I(\mathbf{r}^{\prime})=1$ in $a\leq y\leq a$ and zero outside:
$W(\mathbf{r}_{1},\mathbf{r}_{2})=\frac{1}{4\pi^{2}}\int_{-a}^{a}\frac{e^{-j\beta(R_{2}(y^{\prime})-R_{1}(y^{\prime}))}}{R_{1}(y^{\prime})R_{2}(y^{\prime})}\,dy^{\prime}$ (28)
where $\mathbf{r}_{1}=(z_{1},y_{1})$, $\mathbf{r}_{2}=(z_{2},y_{2})$, $z_{1}>0,z_{2}>0$, and
$R_{n}(y^{\prime})=\sqrt{z_{n}^{2}+(y_{n}-y^{\prime})^{2}}\,\,\,\,\,\,\,\,\,\,n=1,2$ (29)
@@
$W(\mathbf{r}_{1},\mathbf{r}_{2})\approx$ $\frac{a\cdot e^{-j\beta\Phi_{12}}}{z_{1}z_{2}}$
$\cdot\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right)$ (31)
@@
where $a$ is a constant and $e^{-j\beta\Phi_{12}}$ accounts for the difference in average propagation distance from the aperture to each observation point.
@@
$|\mu(\mathbf{r}_{1},\mathbf{r}_{2})|\approx\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right)$ (33)
depends only on the difference in optical path lengths from the edges of the aperture to the observation points [^18] The connection with the Geometrical Theory of Diffraction is a natural consequence of the similarity between coherence integrals and diffraction integrals..
@@
SIFs, therefore, coincide with the curves of constant $\xi$ defined in [18]. Additionally, the mutual information is almost null between any two points belonging to different hyperbolas in the set $R_{2}^{+}-R_{2}^{-}=n\lambda$. This subset of SIFs will be called $Q_{SIF}$ and coincides with the subset $Q_{\xi}$  [^19] As noted in the footnote of Section 3.2, for observation surfaces, the sampling approach yields a slightly higher $N_{DoF}\,\,$. Consequently, the sets $Q_{\xi}$ and $Q_{SIF}$ are not equal. . The equivalence of these two sets, one defined in approximation theory and the other in EIT, is a consequence of the fact that they both represent the same physical limitations in different domains of application.
From a ’spatial information’ perspective, the linear source effectively perceives its surroundings as a hyperbolically warped space  [^20] The warping is given by the conformal transformation $w(z)=\operatorname{arccosh}\left(-{i\zeta}/a\right)$, with the complex variable $\zeta=y+iz$. The approach in [43] implicitly identifies the warping transformation for curved observation domains by working directly on the radiation operator. The identification of warp transformations for surface observation domains remains a challenging and active area of research [44]., as illustrated in Fig. 3. The transformation linearizes the hyperbolas (red curves) and ellipses (blue curves) of Fig. 2. This new geometry provides a clear visualization of how the Degrees of Freedom are distributed among the different receiving domains.
@@
The figure shows that the angular power spectrum in $\mathbf{r}_{1}=\bar{\mathbf{r}}$ is strongly concentrated towards $k_{s}=\beta$. This indicates that the hyperbola is aligned with the optical ray passing through $\bar{\mathbf{r}}$ . For all points to the right of the dashed line, the hyperbola tends toward a straight line, and the $\xi$ curve coincides with the optical ray passing through $\bar{\mathbf{r}}$. In this region, the cross-spectral density (CSD) function locally approximates that of a coherent plane wave, and the hyperbola also coincides with the SIF passing through $\bar{\mathbf{r}}$.

### v3

@@ 4 Spatially incoherent homogeneous linear source
Now, let us consider the EIT stochastic approach. The CSD in Eq. (19) for a spatially uncorrelated, homogeneous line source with spectral density $S_{J}(\mathbf{r}^{\prime})=1$ on the segment $\mathbf{r}^{\prime}=(0,y^{\prime})$, $a\leq y^{\prime}\leq a$, and zero elsewhere, is
$W(\mathbf{r}_{1},\mathbf{r}_{2})\;=\;\frac{(\bar{\omega}\mu_{0})^{2}}{16\pi^{2}}\int_{-a}^{a}\frac{e^{+\,j\beta\,[\,R_{2}(y^{\prime})-R_{1}(y^{\prime})\,]}}{R_{1}(y^{\prime})\,R_{2}(y^{\prime})}\,dy^{\prime},$ (28)
where $\mathbf{r}_{n}=(z_{n},y_{n})$ with $z_{n}>0$ and
$R_{n}(y^{\prime})\;=\;\sqrt{\,z_{n}^{2}+(y_{n}-y^{\prime})^{2}\,},\qquad n=1,2.$ (29)
@@
$W(\mathbf{r}_{1},\mathbf{r}_{2})$ $\;\approx\;\frac{(\bar{\omega}\mu_{0})^{2}}{16\pi^{2}}\,\frac{2a}{z_{1}z_{2}}\;e^{+j\beta\,\Phi_{12}}\;\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right),$ (31)
@@
where $\mathrm{sinc}(x)\equiv\sin x\,/\,x$.
@@
$\big|\mu(\mathbf{r}_{1},\mathbf{r}_{2})\big|\;\approx\;\mathrm{sinc}\left(\beta\frac{(R_{2}^{+}-R_{2}^{-})-(R_{1}^{+}-R_{1}^{-})}{2}\right),$ (33)
depends only on the difference in optical path lengths from the two aperture edges to the observation points. [^18] The connection with the Geometrical Theory of Diffraction follows from the close analogy between coherence (CSD) integrals and classical diffraction integrals.
@@
SIFs, therefore, coincide with the curves of constant $\xi$ defined in [18]. Additionally, the mutual information is almost null between any two points belonging to different hyperbolas in the set $R_{2}^{+}-R_{2}^{-}=n\lambda$. This subset of SIFs will be called $Q_{SIF}$ and coincides  [^19] As noted in the footnote of Section 3.2, on observation surfaces the sampling approach can yield a slightly larger number of samples compared to the $N_{DoF}\,\,$, because achieving a prescribed approximation accuracy typically requires a modest enlargement of the usable bandwidth with respect to the effective bandwidth introduced in [17]. However, if the sampling is chosen consistently with the field’s effective bandwidth value, then $Q_{\xi}$ and $Q_{\mathrm{SIF}}$ coincide in 2D geometry. Loosely speaking, the effective bandwidth marks the boundary between packing and covering descriptions of the received–field set [5]. with the subset $Q_{\xi}$. The equivalence of these two sets, one defined in approximation theory and the other in EIT, is a consequence of the fact that they both represent the same physical limitations in different domains of application.
From a ’spatial information’ perspective, the linear source effectively perceives its surroundings as a hyperbolically warped space  [^20] The warping is given by the conformal transformation $w(z)=\operatorname{arccosh}\left(-{i\zeta}/a\right)$, with the complex variable $\zeta=y+iz$. The approach in [42] implicitly identifies the warping transformation for curved observation domains by working directly on the radiation operator. The identification of warp transformations for surface observation domains remains a challenging and active area of research [43]., as illustrated in Fig. 3. The transformation linearizes the hyperbolas (red curves) and ellipses (blue curves) of Fig. 2. The transformation linearizes the hyperbolas (red curves) and ellipses (blue curves) of Fig. 2. This new geometry provides a clear visualization of how the Degrees of Freedom are distributed across the different receiving domains, effectively modifying the observation-domain geometry perceived by the antenna.
@@
The figure shows that the angular spectrum in $\mathbf{r}_{1}=\bar{\mathbf{r}}$ is strongly concentrated towards $k_{s}=\beta$. This indicates that the hyperbola is aligned with the optical ray passing through $\bar{\mathbf{r}}$ . For all points to the right of the dashed line, the hyperbola tends toward a straight line, and the $\xi$ curve coincides with the optical ray passing through $\bar{\mathbf{r}}$. In this region, the cross-spectral density (CSD) function locally approximates that of a coherent plane wave, and the hyperbola also coincides with the SIF passing through $\bar{\mathbf{r}}$.
@@ 5 Conclusions
Appendix A Appendix: Equivalence of Coherent Modes and Singular Functions for Spatially Incoherent Sources
Notation.
In the following, the symbol $\dagger$ denotes the conjugate transpose (Hermitian transpose) for finite-dimensional quantities such as vectors and dyadics, and the Hilbert-space adjoint for operators with respect to the $L^{2}$ inner product.
Hypothesis
Let $D\subset\mathbb{R}^{3}$ be a bounded, measurable source domain and $S_{\mathrm{obs}}\subset\mathbb{R}^{3}$ a measurable observation domain. Consider a linear, time-invariant (LTI) medium characterized by a dyadic Green’s function $\bar{\bar{\mathbf{G}}}(\mathbf{r},\mathbf{r}^{\prime})\in L^{2}(S_{\mathrm{obs}}\times D;\mathbb{C}^{3\times 3})$.
  1. 1.
    The Radiation Operator: Define the Hilbert–Schmidt (and therefore compact) linear integral operator $\mathcal{A}:L^{2}(D;\mathbb{C}^{3})\to L^{2}(S_{\mathrm{obs}};\mathbb{C}^{3})$ as:
    $(\mathcal{A}\mathbf{J})(\mathbf{r})=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r},\mathbf{r}^{\prime})\mathbf{J}(\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime},\qquad\mathbf{r}\in S_{\mathrm{obs}}.$
    Let $\{\mathbf{u}_{n}\}_{n\geq 1}$ be the set of orthonormal left singular functions and $\{\sigma_{n}\geq 0\}_{n\geq 1}$ the corresponding singular values of $\mathcal{A}$, defined by the eigenvalue problem:
    $\mathcal{A}\mathcal{A}^{\dagger}\,\mathbf{u}_{n}=\sigma_{n}^{2}\,\mathbf{u}_{n}.$
  2. 2.
    The Stochastic Source: Let the source $\mathbf{J}(\mathbf{r}^{\prime})$ be a zero-mean, spatially delta-correlated random vector field within $D$. Its cross-spectral density (CSD) dyadic is
    $\bar{\bar{\mathbf{W}}}_{J}(\mathbf{r}^{\prime}_{1},\mathbf{r}^{\prime}_{2})=\mathbb{E}\big[\mathbf{J}(\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{2})^{\dagger}\big]=\bar{\bar{\mathbf{I}}}\,\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2}),$
    where $\bar{\bar{\mathbf{I}}}$ is the $3\times 3$ identity dyadic and $\mathbb{E}[\cdot]$ denotes statistical expectation.
  3. 3.
    The Radiated Field: The random electric field $\mathbf{E}(\mathbf{r})$ on $S_{\mathrm{obs}}$ results from the linear propagation $\mathbf{E}=\mathcal{A}\mathbf{J}$. Its statistical properties are described by the field CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})=\mathbb{E}\big[\mathbf{E}(\mathbf{r}_{1})\mathbf{E}(\mathbf{r}_{2})^{\dagger}\big]$. This kernel defines a positive semidefinite integral operator $\mathcal{W}$, whose orthonormal eigenfunctions $\{\bm{\psi}_{n}\}$ (the coherent modes) and non-negative eigenvalues $\{\lambda_{n}\}$ are the solutions to the eigenvalue problem:
    $(\mathcal{W}\bm{\psi}_{n})(\mathbf{r}_{1})=\int_{S_{\mathrm{obs}}}\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})\,\bm{\psi}_{n}(\mathbf{r}_{2})\,d\mathbf{r}_{2}=\lambda_{n}\,\bm{\psi}_{n}(\mathbf{r}_{1}).$
Thesis
The set of coherent modes $\{\bm{\psi}_{n}\}$ of the radiated random field is identical (up to a global phase factor) to the set of left singular functions $\{\mathbf{u}_{n}\}$ of the radiation operator $\mathcal{A}$. Their corresponding eigenvalues are equal to the squared singular values:
$\bm{\psi}_{n}(\mathbf{r})=e^{i\phi_{n}}\,\mathbf{u}_{n}(\mathbf{r}),\qquad\lambda_{n}=\sigma_{n}^{2},\qquad\forall\,n\geq 1,$
where $e^{i\phi_{n}}$ represents an arbitrary global phase factor for each mode.
Demonstration
The proof shows that the CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ is identical to the kernel of the operator $\mathcal{A}\mathcal{A}^{\dagger}$.
  1. 1.
    Compute the field CSD kernel $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ by propagating the source CSD:
    $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})$ $=\mathbb{E}\big[(\mathcal{A}\mathbf{J})(\mathbf{r}_{1})(\mathcal{A}\mathbf{J})(\mathbf{r}_{2})^{\dagger}\big]$
    $=\mathbb{E}\!\left[\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{1})\,d\mathbf{r}^{\prime}_{1}\right)\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})\mathbf{J}(\mathbf{r}^{\prime}_{2})\,d\mathbf{r}^{\prime}_{2}\right)^{\dagger}\right]$
    $=\iint_{D\times D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\,\mathbb{E}\!\big[\mathbf{J}(\mathbf{r}^{\prime}_{1})\mathbf{J}(\mathbf{r}^{\prime}_{2})^{\dagger}\big]\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})^{\dagger}\,d\mathbf{r}^{\prime}_{1}\,d\mathbf{r}^{\prime}_{2}$
    $=\iint_{D\times D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime}_{1})\,\big(\bar{\bar{\mathbf{I}}}\,\delta(\mathbf{r}^{\prime}_{1}-\mathbf{r}^{\prime}_{2})\big)\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime}_{2})^{\dagger}\,d\mathbf{r}^{\prime}_{1}\,d\mathbf{r}^{\prime}_{2}$
    $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}.$
  2. 2.
    Find the integral kernel of the operator $\mathcal{A}\mathcal{A}^{\dagger}$. For any $\mathbf{f}\in L^{2}(S_{\mathrm{obs}};\mathbb{C}^{3})$:
    $(\mathcal{A}\mathcal{A}^{\dagger}\mathbf{f})(\mathbf{r}_{1})$ $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,(\mathcal{A}^{\dagger}\mathbf{f})(\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime}$
    $=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\!\left(\int_{S_{\mathrm{obs}}}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\mathbf{f}(\mathbf{r}_{2})\,d\mathbf{r}_{2}\right)d\mathbf{r}^{\prime}$
    $=\int_{S_{\mathrm{obs}}}\left(\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}\right)\mathbf{f}(\mathbf{r}_{2})\,d\mathbf{r}_{2}.$
    By inspection, the kernel of $\mathcal{A}\mathcal{A}^{\dagger}$, let’s call it $\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})$, is
    $\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})=\int_{D}\bar{\bar{\mathbf{G}}}(\mathbf{r}_{1},\mathbf{r}^{\prime})\,\bar{\bar{\mathbf{G}}}(\mathbf{r}_{2},\mathbf{r}^{\prime})^{\dagger}\,d\mathbf{r}^{\prime}.$
  3. 3.
    Conclusion: Comparing the results of steps 1 and 2 shows the kernels are identical, $\bar{\bar{\mathbf{W}}}(\mathbf{r}_{1},\mathbf{r}_{2})\equiv\bar{\bar{\mathbf{K}}}_{\mathbf{u}}(\mathbf{r}_{1},\mathbf{r}_{2})$, which implies the operators themselves are identical:
    $\mathcal{W}\equiv\mathcal{A}\mathcal{A}^{\dagger}.$
    Since the operators are the same, their respective eigenvalue problems, $\mathcal{W}\bm{\psi}_{n}=\lambda_{n}\bm{\psi}_{n}$ and $(\mathcal{A}\mathcal{A}^{\dagger})\mathbf{u}_{n}=\sigma_{n}^{2}\mathbf{u}_{n}$, must be equivalent. It follows that their eigenfunctions, $\bm{\psi}_{n}$ and $\mathbf{u}_{n}$, form the same orthonormal basis (up to arbitrary phase factors), and their eigenvalues are equal, $\lambda_{n}=\sigma_{n}^{2}$. This completes the proof.